A body is pushed giving it on initial velocity of 5m/s and comes to rest after travelling 4.9m . Find the co-officiant of friction? 

Grade:11

1 Answers

Mayank Ranka
2 years ago
Hi,


is the mass of the body is given ?
v2=u2+2as
u=5 m/s and v=0 as the body comes at rest.
so 0=52+(-2a×4.9)
from here we get a =2.55m/s2
by newton’s 2ndlaw, in this case -
fk=m×a. you can get coff. of friction if you know the mass by putting a=2.55m/s2

Thanks
Regards
Ankit Gupta